Mochi:

Jordan plans to add a porch to her house. The new 16’ x 18’ porch will cost an average of $27 per square foot. Jordan also plans to have a 15' sidewalk poured. The cost of the sidewalk will be an additional $35 per foot. The gate on the porch will cost $500.00 extra. What will the total cost be?

Mercury:

three things to consider: cost of the porch, cost of sidewalk, and cost of gate the porch: L*W*(cost per foot) = 16*18*27 the sidewalk: length*cost per foot = 15*35 then the gate: 500 so add these together, be careful about order of operations